OpenStudy (anonymous):

Paul teaches the guitar in groups of three students. He rents a room for the guitar classes at $250 per month. He charges $30 per hour per student. Paul teaches six groups of students, and his net earnings per month are $1,910. Part A: If each student learns guitar for the same number of hours, write an equation that can be used to calculate the number of hours each student learns guitar in a month. Part B: How many hours did each student learn guitar each month? Show your work and justify each step of your work.

OpenStudy (anonymous):

@agent0smith

OpenStudy (anonymous):

@skullpatrol

OpenStudy (anonymous):

30x+250=1910 i do not think what i have is right for part a

OpenStudy (agent0smith):

He rents the room for $250. That means he PAYS 250, he doesn't make 250.

OpenStudy (anonymous):

o ok

OpenStudy (anonymous):

what would be the equation

OpenStudy (agent0smith):

Paul teaches the guitar in groups of three students. Paul teaches six groups of students, How many students must there be, from these two sentences?

OpenStudy (anonymous):

3?

OpenStudy (agent0smith):

Read it again.

OpenStudy (anonymous):

6 groups

OpenStudy (anonymous):

ooh... wait

OpenStudy (anonymous):

18

OpenStudy (agent0smith):

Yes. Now try writing the equation if x is the number of hours a student gets. You had it pretty close above.

OpenStudy (anonymous):

ok.....

OpenStudy (anonymous):

18(30)-250=1910?

OpenStudy (agent0smith):

18 students, 30 an hour per student, with x number of hours means 18*30*x

OpenStudy (agent0smith):

But close!

OpenStudy (anonymous):

18*30*x?

OpenStudy (anonymous):

540x?

OpenStudy (agent0smith):

All you have to do is modify your above equation, you almost had it.

OpenStudy (anonymous):

540x+250=1910

OpenStudy (anonymous):

?

OpenStudy (agent0smith):

But... I thought we went over him paying 250, not making 250

OpenStudy (anonymous):

but dont you need to add it so then you can subtract it when solving?

OpenStudy (anonymous):

u know inverse operations?

OpenStudy (anonymous):

can we go to part b @agent0smith

OpenStudy (agent0smith):

It's nothing to do with inverse operations. He PROFITS 540x, he loses 250, the way you wrote it he's not paying 250 for the room, he's making 250.

OpenStudy (anonymous):

540x-250=1910

OpenStudy (anonymous):

it makes sense

OpenStudy (agent0smith):

Better, now solve it.

OpenStudy (anonymous):

290x=1910

OpenStudy (anonymous):

x=about 6

OpenStudy (anonymous):

?

OpenStudy (agent0smith):

You can't combine unlike terms like that. You can't combine an x term like 540x, with a constant term like -250

OpenStudy (anonymous):

how do you solve then

OpenStudy (anonymous):

@agent0smith

OpenStudy (anonymous):

?

OpenStudy (anonymous):

okay the total answer i got was x=3

OpenStudy (agent0smith):

540x-250=1910 add 250 to both sides

OpenStudy (anonymous):

okay..... you get 540x=2160

OpenStudy (anonymous):

x=4

OpenStudy (anonymous):

i think i am getting the hang of it...... is it right?

OpenStudy (anonymous):

1910 is the NET cost

OpenStudy (anonymous):

that includes 250

OpenStudy (anonymous):

so it would be 540x+250=1910 or 540x-250=1910

OpenStudy (anonymous):

@agent0smith

OpenStudy (agent0smith):

"and his net earnings per month are $1,910"

OpenStudy (agent0smith):

Net earnings are the money he makes overall, after paying for the room.

OpenStudy (anonymous):

yeah so it would be 540x+250=1910

OpenStudy (anonymous):

this is craziness!

OpenStudy (agent0smith):

He MAKES money off his students. He LOSES money to the room, meaning you must subtract it.

OpenStudy (anonymous):

oooooooh he LOSES

OpenStudy (agent0smith):

It's not craziness. We already went over it, you had everything correct.

OpenStudy (anonymous):

now it makes sense i am stupido!